# Hebrew Diacritics Restoration using Visual Representation

### v2

@@ Abstract
In this work, we present DiVRit, a novel system for Hebrew diacritization that frames the task as a zero-shot classification problem. Our approach operates at the word level, selecting the most appropriate diacritization pattern for each undiacritized word from a dynamically generated candidate set, conditioned on the surrounding textual context. A key innovation of DiVRit is its use of a Hebrew Visual Language Model, which processes undiacritized text as an image, allowing diacritic information to be embedded directly within the input’s vector representation.
@@ 1 Introduction
In this research we present DiVRit, a Diacritizer with Visual Representation for Hebrew (IVRIT) text. We explore the use of a vision transformer (ViT) architecture He et al. (2022) as a vision-language model (VLM) to represent diacritized Hebrew text visually. Pretraining on images of rendered undiacritized text establishes a robust visual feature space Rust et al. (2023) essential for modeling diacritics as visual elements. Since undiacritized and diacritized texts share the same visual dimensions, we can further train our model on a diacritized corpus. This allows us to train the VLM to extract distinct embeddings for each candidate Ye et al. (2021) and investigate its potential to differentiate subtle diacritization patterns through visual representations. This differentiation is refined through a scoring mechanism, where the likelihood of each candidate is determined by an inner product with the contextual embedding, a method well-suited for measuring alignment in high-dimensional spaces. The model’s success in selecting the best candidate directly reflects its ability to learn the underlying patterns and logic of diacritization, aligning with the principles of zero-shot learning where the model generalizes to unseen classes based on learned representations.
While our system does not yet surpass Nakdimon’s reported word-level accuracy of 89.75% in all scenarios, initial results demonstrate the promise of our approach. In an “oracle” setting, where the correct diacritization pattern is guaranteed to be within the model’s candidate set, our model achieves word-level accuracy of 92.68%, highlighting its robust ability to assign high scores to correct candidates when they are present. In a more confined setting using a KNN-based candidate generator operating by character sequence alignment, the system reaches 87.87% accuracy. Although slightly below current benchmarks, this still reflects competitive performance. This sets the bar for visual models performing diacritization at the word level operating in a zero-shot learning framework, selecting from full-form candidates rather than predicting individual diacritics. These results suggest that further improvements in candidate generation and training methodologies can significantly enhance its practical effectiveness. Furthermore, DiVRit is the first visual model to perform a candidate-ranking task since these models have been introduced for NLP Salesky et al. (2021). We expect its reliance on visual representations to add diacritics to outperform previous methods especially in the environments identified as strong suits of PIXEL models, such as data noised by typographical errors and adversarial character swapping, raw visual inputs that avoid OCR pipelines, and more. [^1] We will release our data and code upon publication.
@@ 3.1 Architecture
Our processing pipeline integrates a visual candidate encoder, based on the PIXEL-base model Rust et al. (2023), and a context encoder (see Figure 1). This design allows us to process diacritization candidates, which are initially generated for the input undiacritized word (§3.3). Each of these candidates is then rendered as an image and processed by the candidate encoder, generating candidate-specific embeddings that are subsequently used for scoring. Scoring is computed by comparing these candidate embeddings against a contextual embedding of the undiacritized word, extracted by the context encoder which may have visual and/or textual components. Since the encoders may produce multiple embeddings for a single word, either due to their window size limitations in the visual modality or due to subword tokenization in textual encoders , we aggregate such cases into a single vector per item to enable simple scoring via inner product between the contextual word and each candidate embedding. In all cases, we apply mean pooling to ensure a unified representation, preserving information from the entire input span Reimers and Gurevych (2019). To enable direct comparison, both candidate and contextual embeddings are projected to a shared embedding space of identical dimensionality. In our experiments, we explored various architectural configurations to encourage the model to develop a more nuanced and effective representation and prediction space of diacritics.
@@ 3.3 Candidate Generation
While the other diacritizers utilize model embeddings to directly predict or refine diacritization, our approach distinguishes itself by incorporating a candidate generation mechanism. Dicta’s Nakdan utilizes learned representations and human-curated linguistic resources to identify and refine a set of plausible diacritizations, whereas Nakdimon and MenakBERT employ a minimalist, neural-only character-by-character prediction approach that allows them to generate any potential diacritization sequence. In contrast, we first generate candidates for each word using a corpus-driven mechanism, and subsequently employ our architecture to score and select the most contextually-appropriate option from this generated set.
@@ 4.1 Experimental Setup
To overcome these limitations in both candidate and context encoding, we developed a second variant of the model, where the context encoder is instantiated by a text-based AlephBertGimmel-small Hebrew language model Gueta et al. (2023), and the visual candidate encoder is employed following additional pre-training. We found that pretraining the candidate encoder on unmasking diacritized images for an additional 200K steps yielded the best performance gain, presumably because it helps the model learn diacritic patterns and establish more robust diacritic representations. The 200K-step pretraining proved optimal, as increasing it to 400K steps provided a slight degradation of 0.3 percentage points in word-level accuracy. Consequently, for all our main experiments, we chose to utilize this 200K-step pretraining for the candidate encoder, and the context encoder was instantiated by the AlephBertGimmel-small model, which demonstrated superior contextual understanding compared to the vision-based alternative. Following these architectural and pretraining selections, the full DiVRit architecture was trained for an additional 240K steps on the classification objective using multiclass cross-entropy loss, with each step batching 32 examples across two GPUs.
@@ 4.2 Results
Table 1 presents the evaluation results for DiVRit’s best setup, as detailed in Section 4.1, alongside those of the other methods. DiVRit operates well among the ranks of similarly data-driven systems. Moreover, both evaluation variants pass both baselines on all metrics, indicating the potential of visual representations for informing Hebrew diacritics. The Oracle setup greatly outperforms Nakdimon, especially on word-level metrics, and is on par with MenakBERT at the character-level metrics. Surprisingly, despite its reliance on individual character segmentation for in-word representation, MenakBERT excels on word-level metrics in our evaluation. [^2] We note that we did not implement either of the two data-driven character-level prediction systems and are relying on their own reporting. Finally, the gap between the two setups highlights the room left to cover by improving candidate generation, which is our top priority for future work.
@@ 4.3 System Component Analysis
However, longer finetuning does not always guarantee improved performance, so the exact finetuning steps for each setup are detailed in the tables. In this auxiliary task setup, extending the duration to 240K steps significantly decreased accuracy to 82.45% for two-candidate selection and 63.19% for three-candidate selection. This suggests that while auxiliary tasks might guide the model towards certain beneficial representations, they can also impose constraints on the flexibility of the embedding space. Consequently, we shifted our auxiliary task to predict both the diacritics themselves and their corresponding positions within the image embeddings, hypothesizing that this Auxiliary Pos. Encoding setup would embed more detailed information about the diacritic patterns. Its performance for two-candidate selection was 89.17%, a slight decrease compared to both the Basic setup and the Bag-of-Diacritics auxiliary task. However, for three-candidate selection, this setup achieved 76.56% accuracy. This represents a notable improvement in generalization compared to both those setups, demonstrating that a better representation of the diacritics can indeed lead to better generalization of diacritical patterns.
@@ Acknowledgements
DiVRit is a stepping-stone on the road to truly structure-agnostic visual diacritization, but it’s not quite there yet. It still relies on a data-driven candidate generation method, and the best context encoder formation we found is still text-based. Being developed over the Hebrew diacritization task, application to other languages may prove nontrivial, for example given Arabic’s multiple representations for individual characters. This is not an issue for Unicode-based encodings (which treat all forms of a character as the same) and is not a serious issue for visual Hebrew, where only five letters exhibit a mild form of this behavior. Finally, we note that our experiments were performed over a modest setup limited by access to GPUs. Given more resources, we might have been able to achieve better performance with a larger architecture for both encoder models, with a longer pre-training process, and with more runs that allow comprehensive hyperparameter tuning.

### v3

@@ Abstract
In this work, we present DiVRit, a novel system for Hebrew diacritization that frames the task as a zero-shot classification problem. Our approach operates at the word level, selecting the most appropriate diacritization pattern for each undiacritized word from a dynamically generated candidate set, conditioned on the surrounding textual context. A key innovation of DiVRit is its use of a Hebrew Visual Language Model to process diacritized candidates as images, allowing diacritic information to be embedded directly within their vector representations while the surrounding context remains tokenization-based.
@@ 1 Introduction
In this research we present DiVRit, a Diacritizer with Visual Representation for Hebrew (IVRIT) text. We explore the use of a vision transformer (ViT) architecture He et al. (2022) as a vision-language model (VLM) to represent diacritized Hebrew text visually. Pretraining on images of rendered undiacritized text establishes a robust visual feature space Rust et al. (2023) essential for modeling diacritics as visual elements. Since undiacritized and diacritized texts share the same visual dimensions, we can further train our model on a diacritized corpus. This allows us to train the VLM to extract distinct embeddings for each candidate Ye et al. (2021) and investigate its potential to differentiate subtle diacritization patterns through visual representations. The model’s success in selecting the best candidate directly reflects its ability to learn the underlying patterns and logic of diacritization, aligning with the principles of zero-shot learning where the model generalizes to unseen classes based on learned representations.
While our system does not yet surpass Nakdimon’s reported word-level accuracy of 89.75% in all scenarios, initial results demonstrate the promise of our approach. In an “oracle” setting, where the correct diacritization pattern is guaranteed to be within the model’s candidate set, our model achieves word-level accuracy of 92.68%, highlighting its robust ability to assign high scores to correct candidates when they are present. In a more confined setting using a KNN-based candidate generator operating by character sequence alignment, the system reaches 87.87% accuracy. Although slightly below current benchmarks, this still reflects competitive performance. This sets the bar for visual models performing diacritization at the word level operating in a zero-shot learning framework, selecting from full-form candidates rather than predicting individual diacritics. These results suggest that further improvements in candidate generation and training methodologies can significantly enhance its practical effectiveness. Furthermore, DiVRit is the first visual model to perform a candidate-ranking task since these models have been introduced for NLP Salesky et al. (2021). We expect its reliance on visual representations to add diacritics to outperform previous methods especially in the environments identified as strong suits of PIXEL models, such as data noised by typographical errors and adversarial character swapping, raw visual inputs that avoid OCR pipelines, and more. Our code is publicly available online.111github.com/elboheryair/DiVRit
@@ 3.1 Architecture
Our processing pipeline integrates a visual candidate encoder, based on the PIXEL-base model Rust et al. (2023), and a context encoder (see Figure 1). This design allows us to process diacritization candidates, which are initially generated for the input undiacritized word (§3.3). Each of these candidates is then rendered as an image and processed by the candidate encoder, generating candidate-specific embeddings that are subsequently used for scoring. Scoring is computed by comparing these candidate embeddings against a contextual embedding of the undiacritized word, extracted by the context encoder which may have visual and/or textual components. Since the encoders may produce multiple embeddings for a single word, either due to their window size limitations in the visual modality or due to subword tokenization in textual encoders, we aggregate such cases into a single vector per item. This enables simple score calculation as an inner product between the embedding of the context word and each candidate, a method well-suited for measuring alignment in high-dimensional spaces. In all cases, we apply mean pooling to ensure a unified representation, preserving information from the entire input span Reimers and Gurevych (2019). To enable direct comparison, both candidate and contextual embeddings are projected to a shared embedding space of identical dimensionality. In our experiments, we explored various architectural configurations to encourage the model to develop a more nuanced and effective representation and prediction space of diacritics.
@@ 3.3 Candidate Generation
While the other diacritizers utilize model embeddings to directly predict or refine diacritization, our approach distinguishes itself by incorporating a constrained candidate generation mechanism. Nakdimon and MenakBERT employ a minimalist, neural-only character-by-character prediction approach that allows for the generation of any potential sequence. In contrast, Dicta’s Nakdan operates as a hybrid system that leverages human-curated linguistic resources and a formal lexicon. Before ranking, it applies morphological analysis and predicts POS tags to filter irrelevant options from its internal dictionary of pre-annotated word forms. While this ensures high grammatical accuracy, it relies heavily on predefined rules and curated resources. Our system, DiVRit, bridges these two worlds. Similar to Nakdan, we first generate candidates using a corpus-driven mechanism, capturing the empirical distribution of diacritized forms in real-world usage. However, unlike the rule-based constraints of Nakdan’s filtering, our architecture allows the VLM to learn unconstrained representations of these candidates. By processing candidates as images, DiVRit captures diacritic information natively, echoing the flexible, data-driven nature of Nakdimon and MenakBERT.
@@ 4.1 Experimental Setup
To overcome these limitations in both candidate and context encoding, we developed a second variant of the model, where the context encoder is instantiated by a text-based AlephBertGimmel-small Hebrew language model Gueta et al. (2023b), and the visual candidate encoder is employed following additional pre-training. We found that pretraining the candidate encoder on unmasking diacritized images for an additional 200K steps yielded the best performance gain, presumably because it helps the model learn diacritic patterns and establish more robust diacritic representations. The 200K-step pretraining proved optimal, as increasing it to 400K steps provided a slight degradation of 0.3 percentage points in word-level accuracy. Consequently, for all our main experiments, we chose to utilize this 200K-step pretraining for the candidate encoder, and the context encoder was instantiated by the AlephBertGimmel-small model, which demonstrated superior contextual understanding compared to the vision-based alternative. Following these architectural and pretraining selections, the full DiVRit architecture was trained for an additional 240K steps on the classification objective using multiclass cross-entropy loss, with each step batching 32 examples across two GPUs.
@@ 4.2 Results
Table 1 presents the evaluation results for DiVRit’s best setup, as detailed in Section 4.1, alongside those of the other methods. DiVRit operates well among the ranks of similarly data-driven systems. Moreover, both evaluation variants pass both baselines on all metrics, indicating the potential of visual representations for informing Hebrew diacritics. The Oracle setup greatly outperforms Nakdimon, especially on word-level metrics, and is on par with MenakBERT at the character-level metrics. Surprisingly, despite its reliance on individual character segmentation for in-word representation, MenakBERT excels on word-level metrics in our evaluation. [^2] We note that we did not implement either of the two data-driven character-level prediction systems and are relying on their own reporting. Finally, while Dicta understandably sets the performance ceiling, it is a hybrid system that relies on extensive, manually-curated linguistic knowledge. Our work focuses on the more directly comparable data-driven space. In this context, DiVRit’s results are competitive with other neural models like Nakdimon and MenakBERT, particularly in the Oracle setting. This demonstrates that visual representations can effectively internalize complex diacritical patterns natively, offering an interesting and more flexible alternative for scenarios where expert-crafted lexicons or rules are unavailable.
@@ 4.3 System Component Analysis
However, longer finetuning does not always guarantee improved performance, so the exact finetuning steps for each setup are detailed in the tables. In this auxiliary task setup, extending the duration to 240K steps significantly decreased accuracy to 82.45% for two-candidate selection and 63.19% for three-candidate selection. This suggests that while auxiliary tasks might guide the model towards certain beneficial representations, they can also impose constraints on the flexibility of the embedding space. Consequently, we shifted our auxiliary task to predict both the diacritics themselves and their corresponding positions within the image embeddings, hypothesizing that this Auxiliary Pos. Encoding setup would embed more detailed information about the diacritic patterns. Its performance for two-candidate selection was 89.17%, a slight decrease compared to both the Basic setup and the Bag-of-Diacritics auxiliary task. However, for three-candidate selection, this setup achieved 76.56% accuracy. This represents a notable improvement in generalization compared to both those setups, demonstrating that a better representation of the diacritics can indeed lead to better generalization of diacritical patterns. While the positional encoding task is more demanding, limiting absolute peak performance on two-candidate sets, the resulting structural constraint encourages better generalization. This advantage becomes evident as the task complexity grows with the addition of a third candidate.
@@ Acknowledgements
DiVRit is a stepping-stone on the road to truly structure-agnostic visual diacritization, but it’s not quite there yet. It still relies on a data-driven candidate generation method, and the best context encoder formation we found is still text-based. Being developed over the Hebrew diacritization task, application to other languages may prove nontrivial, for example given Arabic’s multiple representations for individual characters. This is not an issue for Unicode-based encodings (which treat all forms of a character as the same) and is not a serious issue for visual Hebrew, where only five letters exhibit a mild form of this behavior. Finally, Our experiments were conducted on a modest GPU setup, so we doubled the training steps to approximate the original PIXEL regimen. With more resources, a larger architecture, longer pre-training, and more extensive hyperparameter tuning could likely yield better results. In particular, Gueta et al. (2023a) show that explicit morphological signals during pre-training substantially improve Hebrew performance on semantic and morphological tasks, implying that hyperparameters such as patch size and masking ratio, which determine how much morphological information must be inferred from partial image patches, may have a particularly strong influence in morphologically rich languages like Hebrew and would benefit from careful optimization. The text rendering strategy is likewise critical Lotz et al. (2023), acting as a visual tokenization step whose impact is especially pronounced in Hebrew due to the diversity of root inflections Gazit et al. (2025). Our current multi-phase curriculum starts with undiacritized text before the diacritized phase. Extending pre-training on the undiacritized base and refining visual hyperparameters could lead to further gains.